OpenStudy (anonymous):

Find the product and quotient functions for f×g and f÷g for the functions given: f(x)=2x+3 and g(x)=-x2+5 f(x)=-x2+15x+8 and g(x)=x2+10x+16 f(x)=2sqrt5−sqrt2 and g(x)=3sqrt5√+2sqrt2

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

these are different and I don't understand them

OpenStudy (anonymous):

i am leaving now for dinner.

OpenStudy (anonymous):

for part the x2 is x^2 sorry

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

f(x)/g(x)=(2sqrt5-sqrt2)/(3sqrt5+2sqrt2)=(2sqrt5-sqrt2)/(3sqrt5+2sqrt2) *(3sqrt5-2sqrt2)/(3sqrt5-2sqrt2) ={6*5-4sqrt10-3sqrt10+2*2}/{9.5-4*2} =(34-7sqrt10)/37

OpenStudy (anonymous):

that was part c? @surjithayer

OpenStudy (anonymous):

part c. f/g

OpenStudy (anonymous):

f(x)*g(x)=(2sqrt5-sqrt2)*(3sqrt5+2sqrt2) =6*5+4sqrt10-3sqrt10-2*2 =26+sqrt10